Two rectangles SO by 50 cm are placed perpendic J arly with a common e Ti 1000 K while surface two balance with 2 large surround ling room at 300 heat lost from the surface at 1000 K to surface number 2.


1
Expert's answer
2021-02-19T07:54:40-0500

Answer

We know Stephen law

E="\sigma T^4"

Where

"\sigma =5.67\times10^{-8}J/T"

Now for given temperature

"E_1=\sigma T_1^4"

"=5.67\times10^{-8}\times(1000) ^4"

"=56700W\/m^2"

"E_2=\sigma T_2^4"

"=5.67\times10^{-8}\times(300) ^4"

"=45027W\/m^2"

Therefore

Temperature of insulated surface

is

"T_2=(\frac{E_2}{\sigma}) ^{0.25}"

Putting the value of E2 and Stephen constant

So

"T_2=(\frac{45027}{5.67\times10^{-8}}) ^{0.25}=566.378K"